The length of a rectangular garden is 1 ft more than 2 times its width. If the perimeter of the garden is 44 ft, find both dimensions of the garden.

Answer by jayanthihemkos@yahoo.com(87)

The length of a rectangular garden is 1 ft more than 2 times its width. If the perimeter of the garden is 44 ft, find both dimensions of the garden.
Let the width of the garden w = x
The length of the garden l = 2x+1
peimeter of a rectangle = 2(l+w)
2(l+w) = 44
2(x+2x+1) = 44
(x+2x+1) = 44/2
3x+1= 22
3x+1-1=22-1
3x = 21
3x/3= 21/3
x = 7ft
l = x = 7 ft
w = 2x+1
w = 2(7) +1
w = 15 ft
length = 7 ft and width = 15 ft is the solution